OpenStudy (anonymous):

When simplified, (1+1/x)/{(1/x^2)-1} equals?

OpenStudy (anonymous):

1/x^3 +1/x^2 -1/x -1

OpenStudy (anonymous):

\[\frac{ 1+\frac{ 1 }{ x } }{ \frac{ 1 }{ x^2 }-1 }\]

OpenStudy (anonymous):

I would make the numerator and denominator into 1 fraction. This is the same as in basic math. find a common denominator and add

OpenStudy (anonymous):

This one is multiple choice. The answers are between a) x/1-x b) x/x-1 c) 1/x-1 d) x-1/x

OpenStudy (anonymous):

\[\frac{ \frac{ x }{ x } +\frac{ 1 }{ x }}{ \frac{ 1 }{ x^2 } -\frac{ x^2 }{ x^2 }}=\frac{ \frac{ x+1 }{ x } }{ \frac{ 1-x^2 }{ x^2 } }\]

OpenStudy (anonymous):

Now just like in basic math when we divide fractions we can flip the denominator and multiply instead

OpenStudy (anonymous):

\[\frac{ x+1 }{ x }\times \frac{ x^2 }{ 1-x^2 }=\frac{ (x+1)x }{ 1-x^2 }\]You can see that an x will cancel from the bottom.

OpenStudy (anonymous):

we can factor the denominator next

OpenStudy (anonymous):

Do you want to learn how to do it or just have the answer? Think wisely before you answer this question

OpenStudy (anonymous):

Sure, could you teach me

OpenStudy (anonymous):

\[\frac{ (x+1)x }{ 1-x^2 }=\frac{ (x+1)x }{ (1-x)(1+x) }=\frac{ x }{ 1-x }\]

OpenStudy (anonymous):

just write down each of my posted steps and ask me any questions you don't understand. I have posted all of my work and I tried to post my thought process as well.

OpenStudy (anonymous):

After studying it a bit, I understand it. Thank you very much.

OpenStudy (anonymous):

yw